# Timescales diagnostics for saving viscous and MHD-driven dusty discs from external photoevaporation

### v1

@@ Appendix B Dust entrainment in MHD winds and role of radial drift
Figure 8: Evolution of the dust components in models including (thick dashed lines) or not including (thin continuous lines) dust entrainment and removal in MHD winds. The top panel represents the subset of the disc parameters used in this paper which would be most affected by MHD-wind dust entrainment, namely $\psi_{\mathrm{DW}}\gg 1$, high $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{DW}}=0.01$, low ${R_{\mathrm{c}}}=10\,\mathrm{AU}$, our fiducial $\lambda=3$, and various FUV fluxes. We plot the evolution of the total dust mass, and the evolution of the dust mass lost to the photoevaporative wind, to the MHD wind and to the star (cfr. second row of Fig. 3). A top inset with a linear vertical axis scaling is added to better display the very small differences which would otherwise be undetectable. The bottom panel shows the same disc setups but with $\lambda=1.5$, where MHD-wind-driven mass removal is most effective. Again, even in this extreme case, adding MHD-wind dust entrainment modifies only slightly the dust evolution and only affects the amount of dust accreted onto the star by a marginal amount (only visible in the top inset with linear vertical axis scaling).

### v2

@@ Appendix B Dust entrainment in MHD winds and role of radial drift
Figure 8: Evolution of the dust components in models including (thick dashed lines) or not including (thin continuous lines) dust entrainment and removal in MHD winds. The top panel represents the subset of the disc parameters used in this paper which would be most affected by MHD-wind dust entrainment, namely $\psi_{\mathrm{DW}}\gg 1$ (MHD-wind discs), high $\alpha_{\mathrm{tot}}=\alpha_{\mathrm{DW}}=0.01$, low ${R_{\mathrm{c}}}=10\,\mathrm{AU}$, our fiducial $\lambda=3$, and various FUV fluxes. We plot the evolution of the total dust mass, and the evolution of the dust mass lost to the photoevaporative wind, to the MHD wind and to the star (cfr. second row of Fig. 3). A top inset with a linear vertical axis scaling is added to better display the very small differences which would otherwise be undetectable. The bottom panel shows the same disc setups but with $\lambda=1.5$, where MHD-wind-driven mass removal is most effective. Again, even in this extreme case, adding MHD-wind dust entrainment modifies only slightly the dust evolution and only affects the amount of dust accreted onto the star by a marginal amount (only visible in the top inset with linear vertical axis scaling).